OpenStudy (anonymous):

How many coefficients are there in the expression: 2c + 6 - 9d – 4.4 – 8d + 3d 1 2 3 4

OpenStudy (anonymous):

Group all like terms with the Distributive Property

OpenStudy (anonymous):

9d + 4d = (9+4) d = 13 d

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

Group all of the c's, d's and no-coefficients together first.

OpenStudy (anonymous):

i dont understand